OpenStudy (cherie_magee):

Given the system of equations, what is the solution? x + 2y = 11 x - 2y = -1 {(-5, -3)} {(1, 1)} {(5, 3)}

OpenStudy (austinl):

We can use elminiation. -1(x + 2y = 11) x - 2y = -1 -x-2y=-11 x-2y=-1 -4y=-12 y=3 Then you plug that back into one of the original equations and solve for x.

OpenStudy (austinl):

Hello?

jhonyy9 (jhonyy9):

x + 2y = 11 x - 2y = -1 ------------ assum the same terms 2x 0 = 10 2x=10 x=10/2 = 5 x=5 5-2y=-1 5+1=2y 6=2y y=6/2 y= 3 x=5 y=3